A speedboat starts from rest and accelerates at +2.01 m/s2 for 7.50 s. At the end of this time, the boat continues for an additional 5.75 s with an acceleration of +0.518 m/s2. Following this, the boat accelerates at −1.49 m/s2 for 8.35 s.
Answer #1

Here,

a) for the velocity at 21.6 s

v = u + at

v = 0 + 2.01 * 7.5 + 0.518 * 5.75 - 1.49 * 8.35

v = 5.612 m/s

the velocity at 21.6 s is 5.61 m/s

b) Now, for the total displacement

total displacement = 0.50 * 2.01 * 7.5^2 + (2.01 * 7.5 * 5.75 + 0.50 * 0.518 * 5.75^2) + ((2.01 * 7.5 + 5.75 * 0.518) * 8.35 - 0.50 * 1.49 * 8.35^2)

total displacement = 150.6 m

the total displacement is 150.6 m